Suppose a company raised $3,000,000 to fund its expansion. It expects the sustainable growth to be 3% a year.

It sold 100 20-year corporate bonds with a $10,000 par value and a 4% coupon rate at the market price of $10,200. The flotation cost is 2% of par value.

The corporate income tax rate is 25%.

It issued 100,000 new preferred shares at $10.30/share. It promises a $0.60/share annual dividend. The flotation cost is $0.30/share.

The company also issued 20,000 new common shares at $51/share. The flotation cost is $1/share. Its next dividend is $2.5/share.

Before-tax cost of capital from bond =   %

After-tax cost of capital from bond (rd) =   %

Cost of capital from preferred stock (rp) =   %

Cost of capital from common stock (rn) =   %

The weighted average cost of capital is (ra) =   %.
Answer #1
Before-tax cost of bonds (assuming annual coupons):
Current market price of the bond= Present value of all its future coupons+PV of Face value to be recd. At maturity
ie.Price/PV= (Coupon amt.*(1-(1+r)^-n)/r)+(FV/(1+r)^n)
where,
PV= Price-Flotation cost,ie. 10200-(2%*10000)= 10000
Coupon amt.= periodic $ interest amt.= 10000*4%= $ 400
r= the yield /market interest rate/cost to be found out ??
n= no.of coupon periods still pending to maturity,ie. 20
So,plugging in the known values in the above formula,
10000= (400*(1-(1+r)^-20)/r)+(10000/(1+r)^20)
Solving for r, we get the before- tax r as
4%
The corporate income tax rate is 25%.
So, the after-tax cost (rd)=
Before-tax cost*(1-Tax rate)
ie. 4%*(1-25%)=
3%
Cost of preferred stock(r p)
$ Dividend/(net proceeds of issue)
ie. $ Dividend/(Price-Flotation cost)
ie. 0.60/(10.30-0.30)
6%
Cost of Common stock
As per Constant dividend growth model,
r n=(Next dividend/(Market price-Flotation cost))+growth rate
ie.(2.5/(51-1))+3%=
8%
WACC=(Wt.d*rd)+(Wt.p*rp)+(Wt.n*rn)
Type of capital Market Value Wt. to total Cost Wt.*cost
Bonds 100*10200= 1020000 33.22% 3% 1.00%
Preferred shares 100000*10.30= 1030000 33.55% 6% 2.01%
Common stock 20000*51= 1020000 33.22% 8% 2.66%
Total 3070000 100.00% 5.67%
WACC =(33.22%*3%)+(33.55%*6%)+(33.22%*8%)=
5.67%
ANSWERS:
Before-tax cost of capital from bond = 4  %
After-tax cost of capital from bond (rd) = 3 %
Cost of capital from preferred stock (rp) = 6  %
Cost of capital from common stock (rn) = 8 %
The weighted average cost of capital is (ra) = 5.67 %.